OpenStudy (anonymous):

2sqrt√208 ____________ 4^3sqrt√135

OpenStudy (anonymous):

1682.87

OpenStudy (anonymous):

no decimals just simplify please

OpenStudy (anonymous):

1683

OpenStudy (anonymous):

\[\frac{2 \sqrt{208}}{4^3\sqrt{135}}\text{=}\frac{2 \left(4 \sqrt{13}\right)}{64\left(3 \sqrt{15}\right)}=\frac{8}{192}\sqrt{\frac{13}{15}}=\frac{1}{24}\sqrt{\frac{13}{15}}\]